OpenStudy (anonymous):

How can the equation x + 2 = 5 be solved using algebra tiles? A. Add two 1–tiles to each side. B. Remove two 1–tiles from the left side. C. Remove two 1–tiles from each side. D. Remove five 1–tiles from each side.

OpenStudy (anonymous):

X + 2 = 5 If we try to add 2 on each side, we get X + 2 + 2 = 5 + 2 which comes out as X + 4 = 7, No help. If we remove 2 from the left side, we get X + 2 -2 = 5 and we'd get X = 5, but if we cannot remove on one side without removing from the other. Algebra rules. If we remove 5 from both sides: X + 2 - 5 = 5 - 5, which is X + (-3) = 0 And we would need another step to determine X So if we remove 2 from both sides, we'd get X + 2 -2 = 5 - 2 which is X = 3, which is a good result. Therefore the answer is C